OpenStudy (anonymous):

I need help guys with the elimination method for 4y+2x=18 3x+6y=26

OpenStudy (jdoe0001):

\(\bf \begin{array}{llll} 4y+2x=18\\ 3x+6y=26 \end{array}\implies \begin{array}{llll} 2x+4y=18\\ 3x+6y=26 \end{array}\impliedby \textit{arranging the variables} \\ \quad \\ \begin{array}{llll} 2x+4y=18&{\color{brown}{ \times 3}}\\ 3x+6y=26&{\color{brown}{ \times -2}} \end{array}\) if you multiply those expressions, by THAT MUCH, what do they become?

OpenStudy (jdoe0001):

that is, you mutiply the top by 3 and the bottom one by -2

OpenStudy (anonymous):

what if I add both equations down like.. 2x+4y=18 +3x+6y=26 equals 5x+10y=44? is this a right step?

OpenStudy (anonymous):

Is this an option?